OpenStudy (anonymous):

If a ball is thrown straight up into the air with an initial velocity of 64 ft/sec, it's height in feet after t seconds is given by y=65t-16t^2. Find the average velocity for the time period beginning at t = 1 and lasting for .5, .1, and .01. I've done this about 10 times and consistently get -40 for the first which lasts .5 seconds which I'm told is incorrect. Help!

OpenStudy (anonymous):

this is quite annoying, but you have to replace t by 1.5,1.1,1.01 and see what you get which one would you like to do?

OpenStudy (anonymous):

lets do 1.1 \[\frac{65\times 1.1-16(1.1)^2-49}{1.1-1}\] \[\frac{52.14-49}{.1}\] \[\frac{3.14}{.1}\] \[31.4\]

OpenStudy (anonymous):

others are similar, use a calculator

OpenStudy (anonymous):

or take this template, and replace 1.1 by your other two values (everywhere you see it) http://www.wolframalpha.com/input/?i=%2865%281.1%29-%2816%29%281.1%29^2-49%29%2F%281.1-1%29

OpenStudy (anonymous):

Thank you muchly!

OpenStudy (anonymous):

yw